Ed and Hew are product testers at IL Incorp. Ed can drop a piece of luggage off the top of headquarters and then record the results once every 4 minutes. Hew can do the same once every 5 minutes. How many minutes will it take them together to drop their luggage and record the results a (combined)total of 18 times?
Found 2 solutions by Boreal, greenestamps:
Answer by Boreal(15235)

in x minutes they do x/5 and x/4 of these
Need an integer answer
(x/5)+(x/4)=18
multiply through by 20
4x+5x=360
9x=360
x=40 minutes

Answer by greenestamps(13334)



Consider the least common multiple of the two given times, which is 20 minutes.

In 20 minutes, Ed can drop 20/4=5 pieces of luggage and Hew can drop 20/5=4 pieces of luggage. So in 20 minutes together, the two of them can drop 5+4=9 pieces of luggage.

18 pieces is twice that number; so it will take them twice as long -- 40 minutes -- to drop 18 pieces of luggage.